a fast plane making a 2400 mile trip was delayed 14 minutes before take off because of bad weather. To make up for the delay, the pilot increased his airspeed and took a favorable tailwind. Thus he was able to average 6 miles per hour more than usual. Find the usual average ground speed and flying time for the trip.
I know the first equation is (x+6)*(y+(14/60))=2400 but i can't figure out the second equation to do a system of equations.
(x=average airspeed y=flying time)

Answer by mananth(16946)

average speed =x
time = y
..
xy=2400
x=2400/y
(x+6)(y-14/60)= 2400
(x+6)(60y-14)/60=2400
(x+6)(60y-14)=60*2400
plug x
(2400/y +6)(60y-14)=60*2400
(2400+6y)(60y-14)/y=144000
(2400+6y)(60y-14)=144000y
6(400+y)*2(30y-7)=144000y
(400+y)(30y-7)=144000y/12
12000y-2800+30y^2-7y=12000y
30y^2-7y-2800=0
solve using quadratic formula
b^2-4ac=336049
...
y=9.78 hours
x=2400/9.78 = 245.4 mph